Please solve and show all work. Thank you!
The sum of a number and its square is 12. Find the number.

Found 2 solutions by wuwei96815, venugopalramana:
Answer by wuwei96815(245)

x^2 + x = 12
x^2 + x - 12 = 0
(x + 4)(x - 3) = 0

The values of x are -4 and 3.

Does that look right?

Answer by venugopalramana(3286)

Please solve and show all work. Thank you!
The sum of a number and its square is 12. Find the number.
X+X^2=12
X^2+X-12=0
X^2+4X-3X-12=0
X(X+4)-3(X+4)=0
(X-3)(X+4)=0
X-3=0....X=3...........OR......X+4=0...THAT IS X=-4
HENCE ANSWER IS X=3 OR -4